# Design and Analysis of Binaural Signal Matching with Arbitrary Microphone Arrays

### v2

@@ 1 Introduction
Binaural reproduction is an ongoing research topic with an increasing number of applications for augmented and virtual reality, teleconferencing and hearing aids [1]. To binaurally reproduce an acoustic scene, the sound field and the head-related transfer functions (HRTFs) are required. In real-life acoustic scenes, these can be captured simultaneously using microphones that are positioned in the ears of a listener or an anatomically equivalent dummy. However, for more flexible reproduction that includes HRTF personalization, head-tracking and spatial manipulation of the sound field, a more complex array of microphones may be required. Therefore, binaural reproduction with such arrays has become a topic of interest recently.
A common approach is to use higher order Ambisonics (HOA) signals for binaural reproduction [2, 3]. Systems that capture and reproduce HOA were proposed in [4] using spatial coding, in [5] using spherical microphone arrays and in [6] for reproduction via loudspeaker arrays. While HOA is a well studied format, it often requires arrays with high directional resolution and a specific configuration, such as spherical arrays. For non-spherical arrays, [7] proposed a solution for elevation-invariant sound-fields with fully-circular arrays. It was later extended in [8] to more general array-geometries of circumferential contours around non-spherical bodies using a numerical least-squares fitting. However, the study included a relatively large array, consisting of 18 microphones. Another work [9] studied Ambisonics reproduction with smaller arrays, but it is limited to a dedicated microphone array with a “three-plus-one” configuration. To summarize, Ambisonics-based reproduction is currently missing a clear framework for smaller arrays, such as mobile, handheld or wearable arrays, that are often desired for many applications.
@@
A third approach for binaural reproduction, which is also flexible in the array geometry, is binaural signals matching (BSM). This usually refers to the estimation of binaural signals while minimizing the mean-squared error (MSE) by matching the array steering vectors to the HRTFs using a linear formulation. Examples for such methods include [24], which optimized the microphone positions in a planar array geometry. It was later extended to include HRTF smoothing for the purpose of producing perceptually accurate binaural signals with a smaller number of microphones [25], and to include different regularization techniques for the MSE minimization [26, 27]. Furthermore, the method was evaluated with a perceptual study that showed it can be used with individual HRTFs and outperform generic HRTFs, but mainly for directions which were directly optimized in the derivations of the filter coefficients [28]. While these works proved that BSM can produce high quality binaural signals, they only studied relatively complex microphone arrays, consisting of 24 high-quality sensors in a planar geometry. This limits the use of BSM with more general array geometries. Another work which studied a least-squares-based reproduction and integrated a magnitude least-squares (MagLS) solution is described in [29], but the study was limited to Ambisonics signals. This work was extended in [30] by integrating the array model into the MagLS objective function. While the methods described in [30] can be used with any array geometry, the developments and evaluations are limited to spherical and equatorial microphone arrays that are suitable to perform a decomposition of at least a subset of SHs. Furthermore, evaluations only included the Eigenmike array [31] and a fully-circular nine-microphone array.
To summarize, current binaural reproduction solutions are limited. Some require relatively complex arrays, i.e., Ambisonics processing and some of the BFBR methods. Others rely on the estimation accuracy of signal-dependent model parameters. Lastly, many solutions are suitable to specific array designs and acoustic environments. While all present useful results, as has been further reviewed in [1], they do not offer a general design framework that can be successfully applied to any array geometry.
In this paper, a theoretical framework for the design of BSM systems is developed. This framework can be utilized with any array geometry, albeit the quality of the reproduced signal may depend on the specific geometry. First, BSM is developed theoretically for sound fields that are comprised of a known number of uncorrelated sound sources with known directions. Then, well-defined conditions that extend the method to other more general sound field types, which are comprised of an arbitrary number of sources and their directions, are developed. Following that, an extension to BSM that improves performance at high frequencies is developed based on MagLS, and a solution for compensating head rotations is described. Next, an objective analysis studying BSM accuracy using MSE and perceptually-motivated measures with a semi-circular array is presented. Finally, a subjective analysis that includes a four-microphone array that is mounted on a pair of glasses is presented using a listening test to validate the theoretical results.
@@
    A perceptually-motivated extension to BSM is developed and tested for high frequencies (where reproduction accuracy and/or quality with arbitrarily small arrays may degrade).
  4. 4.
    A comprehensive simulation study and a listening experiment validate the accuracy and quality of BSM with small microphone arrays. The study includes head rotations, which further motivates the incorporation of BSM in wearable and mobile devices.
@@ 2 Background
This section presented array processing models and binaural reproduction methods. Throughout the paper, the spherical coordinates system will be used, denoted $(r,\theta,\phi)$, where $r$ is the distance from the origin, $\theta$ is the polar angle measured downward from the Cartesian $z$ axis to the $xy$ plane, and $\phi$ is the azimuthal angle measured from the positive $x$ axis towards the positive $y$ axis.
@@ 2.1 Microphone Array Measurement Model
where $\mathbf{x}(k)=\begin{bmatrix}x_{1}(k),x_{2}(k),\ldots,x_{M}(k)\end{bmatrix}^{T}$ is a vector of length $M$ containing the microphone signals, $\mathbf{V}(k)=\begin{bmatrix}\mathbf{v}_{1}(k),\mathbf{v}_{2}(k),\ldots,\mathbf{v}_{Q}(k)\end{bmatrix}$ is an $M\times Q$ matrix with its $q$-th column containing the steering vector of the $q$-th source [17] for $q=1,2,\ldots,Q$, $\mathbf{s}(k)=\begin{bmatrix}s_{1}(k),s_{2}(k),\ldots,s_{Q}(k)\end{bmatrix}^{T}$ is a vector of length $Q$ containing the source signals, $\mathbf{n}(k)=\begin{bmatrix}n_{1}(k),n_{2}(k),\ldots,n_{M}(k)\end{bmatrix}^{T}$ is an additive noise vector of length $M$, and $(\cdot)^{T}$ is the transpose operator. The set of steering vectors,
where $\mathbf{d}_{m}$ is the Cartesian coordinates of the $m$-th microphone in the array and $v(k,\mathbf{d}_{m};\theta_{q},\phi_{q})$ is the transfer function between a far-field source with a DOA of $(\theta_{q},\phi_{q})$ and $\mathbf{d}_{m}$ for $m=1,2,\ldots,M$, can be calculated analytically for various array types [17, 32], numerically, or measured.
@@ 2.2 Binaural Signal Representation using HRTFs
Assume that a listener is surrounded by a sound field that can be described by the plane-wave amplitude density (PWD) function $a(k,\theta,\phi)$. Then, the sound pressure at the listener’s ears can be described by:
@@
is a vector of length $Q$ containing the HRTFs corresponding to the $Q$ directions of the sources. The narrow-band signals presented here can be transformed to the time domain, where they can be played back over headphones for binaural reproduction.
@@ 2.3 Beamforming-based Binaural Reproduction
where $y(k,\theta_{d},\phi_{d})$ is the beamformer output,
$\mathbf{w}_{d}(k)=\begin{bmatrix}w_{1}(\theta_{d},\phi_{d}),w_{2}(\theta_{d},\phi_{d}),\ldots,w_{M}(\theta_{d},\phi_{d})\end{bmatrix}^{T}$ (6)
@@
where $\{\alpha_{d}\}_{d=1}^{D}$ are the scaling factors. In BFBR design, the beamformer weights, the set of look-directions and the scaling factors are chosen with the aim of producing the desired binaural signals. For instance, with spherical arrays they can be set to produce binaural signals that are equivalent to HOA reproduction [23].
@@ 2.4 Binaural Reproduction using Ambisonics
where $\tilde{a}_{nm}(k)$ and $h_{nm}^{l,r}(k)$ are the spherical Fourier transform (SFT) coefficients of the complex conjugate of $a(k,\theta,\phi)$ and of $h^{l,r}(k,\theta,\phi)$, respectively, $N_{p}=\min\{N_{a},N_{H}\}$ and $(\cdot)^{*}$ is the complex conjugate operator. Equation (8) is known as the Ambisonics format for binaural reproduction [2], and it is presented here since this formulation will be used in the listening experiment as a reference.
@@ 3 Proposed Method for Binaural Signal Matching
This section describes the proposed BSM method for binaural reproduction with general arrays. This method produces time-invariant multiple-input-multiple-output (MIMO) filters in the frequency domain. While the derivations of BSM shown here were already presented in previous works [24, 33, 13, 30], they are repeated here for completeness and to highlight the specific design parameters of BSM that are analyzed in the following section. These parameters are crucial for generating accurate binaural signals. The wave-number index, $k$, will be omitted for brevity, but the following derivations remain narrow-band.
@@ 3.1 Formulation of the BSM Approach
$z^{l,r}=[\mathbf{c}^{l,r}]^{H}\mathbf{x},$ (9)
where $\mathbf{c}^{l,r}$ are vectors of length $M$ holding the filter coefficients for the left and rights ears. Next, the following MSE between the binaural signals in (4) and the filtered microphone signals in (9) is calculated for each ear separately:
$\epsilon^{l,r}=\mathbb{E}\begin{bmatrix}|p^{l,r}-z^{l,r}|^{2}\end{bmatrix},$ (10)
where $\mathbb{E}\begin{bmatrix}\cdot\end{bmatrix}$ is the expectation operator. Assuming that the source signals, $\{s_{q}\}_{q=1}^{Q}$, are uncorrelated with the noise components, $\{n_{m}\}_{m=1}^{M}$, and substituting (4) and (9) in (10), leads to the following MSE:
$\epsilon^{l,r}=$ $\big{(}[\mathbf{c}^{l,r}]^{H}\mathbf{V}-[\mathbf{h}^{l,r}]^{T}\big{)}\mathbf{R_{s}}\big{(}[\mathbf{c}^{l,r}]^{H}\mathbf{V}-[\mathbf{h}^{l,r}]^{T}\big{)}^{H}$
@@
where $\mathbf{R_{s}}=\mathbb{E}\begin{bmatrix}\mathbf{ss}^{H}\end{bmatrix}$ and $\mathbf{R_{n}}=\mathbb{E}\begin{bmatrix}\mathbf{nn}^{H}\end{bmatrix}$. When the sound sources and noise are spatially white, these matrices reduce to $\mathbf{R_{s}}=\sigma_{s}^{2}\mathbf{I}_{L}$ and $\mathbf{R_{n}}=\sigma_{n}^{2}\mathbf{I}_{M}$, where $\sigma_{s}^{2}$ and $\sigma_{n}^{2}$ are the source and noise powers, respectively, and $\mathbf{I}_{L}$ and $\mathbf{I}_{M}$ are the identity matrices of sizes $L$ and $M$, respectively. In this case, (11) can be further simplified to
$\epsilon^{l,r}=\sigma_{s}^{2}\left\lVert\mathbf{V}^{T}[\mathbf{c}^{l,r}]^{*}-\mathbf{h}^{l,r}\right\rVert_{2}^{2}+\sigma_{n}^{2}\left\lVert[\mathbf{c}^{l,r}]^{*}\right\rVert_{2}^{2},$ (12)
where $\left\lVert\cdot\right\rVert_{2}$ is the $l^{2}$-norm, and the second term in (12) can be interpreted as Tikhonov-regularization [34]. The following derivations of BSM utilize the simplified model in (12), which requires less information on the sound field compared to the model in (11). Furthermore, it will be shown later that using (12) may not limit the use of BSM to the more general sound fields corresponding to (11).
@@
$\hat{p}_{\text{BSM}}^{l,r}=[\mathbf{c}^{l,r}_{\text{BSM}}]^{H}\mathbf{x},$ (15)
where $\hat{p}_{\text{BSM}}^{l,r}$ are the estimated binaural signals according to the BSM method.
@@
To illustrate the use of BSM for binaural reproduction, Fig. 1 shows a person recording an acoustic environment using a wearable array on the left. The recorded signals are than processed according to (15) which produces the binaural signals. These are then played back over headphones to a remote person, shown on the right.

### v3

@@ 1 Introduction
Binaural reproduction is an ongoing research topic with an increasing number of applications for augmented and virtual reality, teleconferencing and hearing aids [1]. To binaurally reproduce an acoustic scene, the sound field and the head-related transfer functions (HRTFs) are required. In real-life acoustic scenes, these can be captured simultaneously using microphones that are positioned in the ears of a listener or an anatomically equivalent dummy. However, this approach limits the separation of the sound field and HRTF components, impeding the reproduction of the sound field for different head orientations, or using a different HRTF for personalization. To overcome this limitation, an array of microphones is typically required. Therefore, binaural reproduction with microphone arrays has become a topic of interest recently. An illustration of headphone binaural reproduction is presented in Fig. 1, showing a person recording sound from the environment using a wearable array on the left. The recorded signals are then processed to produce binaural signals. These are then played back over headphones to a remote person, shown on the right.
A common approach to decouple the sound field from the effect of the head is to use higher order Ambisonics (HOA) signals for binaural reproduction [2, 3]. Systems that capture and reproduce HOA were proposed in [4] using spatial coding, in [5] using spherical microphone arrays and in [6] for reproduction via loudspeaker arrays. While HOA is a well studied format, it often requires arrays with high directional resolution and a specific configuration, such as spherical arrays. For non-spherical arrays, [7] proposed a solution for elevation-invariant sound-fields with fully-circular arrays. It was later extended in [8] to more general array-geometries of circumferential contours around non-spherical bodies using a numerical least-squares fitting. However, the study included a relatively large array, consisting of 18 microphones. Another work [9] studied Ambisonics reproduction with smaller arrays, but it is limited to a dedicated microphone array with a “three-plus-one” configuration. To summarize, Ambisonics-based reproduction is currently missing a clear framework for smaller arrays and arbitrary microphone positions, such as mobile, handheld or wearable arrays, that are often desired for many applications.
@@
A third approach for binaural reproduction, which is also flexible in the array geometry, is binaural signals matching (BSM). This usually refers to the estimation of binaural signals while minimizing the mean-squared error (MSE) by matching the array steering vectors to the HRTFs using a linear formulation. Examples for such methods include [24], which optimized the microphone positions in a planar array geometry. It was later extended to include HRTF smoothing for the purpose of producing perceptually accurate binaural signals with a smaller number of microphones [25], and to include different regularization techniques for the MSE minimization [26, 27]. Furthermore, the method was evaluated with a perceptual study that showed it can be used with individual HRTFs and outperform generic HRTFs, but mainly for directions which were directly optimized in the derivations of the filter coefficients [28]. While these works proved that BSM can produce high quality binaural signals, they only studied relatively complex microphone arrays, consisting of 24 high-quality sensors in a planar geometry. This limits the use of BSM with more general array geometries. Another work which studied a least-squares-based reproduction and integrated a magnitude least-squares (MagLS) solution is described in [29], but the study was limited to Ambisonics signals. This work was extended in [30] by integrating the array model into the MagLS objective function, thus providing an end-to-end MagLS (eMagLS) binaural renderer. While the methods described in [30] can be used with any array geometry, the developments and evaluations are limited to spherical and equatorial microphone arrays. Implementing these arrays in practice can be relatively complex, requiring specific designs that may be impractical for arrays mounted on small devices. Thus, later works studied methods with smaller arrays having only a few microphones, such as head-worn arrays [31, 32, 33]. These works highlight the superiority of matching to binaural signals, also denoted as eMagLs, for such arrays, compared to other popular non-parametric methods, such as Ambisonics rendering and BFBR. However, without a rigorous theoretical analysis, the conditions under which matching to binaural signals guarantees accurate binaural reproduction remain unclear, limiting its wider applicability and optimization.
In this paper, a theoretical framework for the design of BSM systems, including eMagLS, is developed. The framework provides a theoretical basis for analyzing the performance of BSM reproduction with any array geometry. Following the theoretical analysis of BSM, the method is numerically studied using MSE and perceptually-motivated measures with a semi-circular array. Finally, a listening experiment that includes a four-microphone array that is mounted on a pair of glasses is presented to validate the theoretical results.
@@
    A comprehensive simulation study and a listening experiment validate the accuracy and quality of BSM with small microphone arrays. The study includes head rotations and MagLS, which further motivates the incorporation of BSM in wearable and mobile devices.
@@ 2 Background
This section presented array processing models and binaural reproduction methods. Throughout the paper, a spherical coordinates system will be used, denoted $(r,\theta,\phi)$, where $r$ is the distance from the origin, $\theta$ is the polar angle measured downward from the Cartesian $z$ axis to the $xy$ plane, and $\phi$ is the azimuthal angle measured from the positive $x$ axis towards the positive $y$ axis. This coordinate system including the corresponding reference head orientation that will be used in this study is illustrated in Fig. 2.
Figure 2: An illustration of the spherical coordinate system and the reference head orientation that is used in this study.
@@ 2.1 Microphone Array Measurement Model
where $\mathbf{x}(k)=\begin{bmatrix}x_{1}(k),x_{2}(k),\ldots,x_{M}(k)\end{bmatrix}^{T}$ is a vector of length $M$ containing the microphone signals, $\mathbf{V}(k)=\begin{bmatrix}\mathbf{v}_{1}(k),\mathbf{v}_{2}(k),\ldots,\mathbf{v}_{Q}(k)\end{bmatrix}$ is an $M\times Q$ matrix with its $q$-th column containing the steering vector of the $q$-th source for $q=1,2,\ldots,Q$, $\mathbf{s}(k)=\begin{bmatrix}s_{1}(k),s_{2}(k),\ldots,s_{Q}(k)\end{bmatrix}^{T}$ is a vector of length $Q$ containing the source signals, $\mathbf{n}(k)=\begin{bmatrix}n_{1}(k),n_{2}(k),\ldots,n_{M}(k)\end{bmatrix}^{T}$ is an additive noise vector of length $M$, and $(\cdot)^{T}$ is the transpose operator. The set of steering vectors can be described as
$\mathbf{v}_{q}(k)=\begin{bmatrix}v(k,\mathbf{d}_{1};\theta_{q},\phi_{q})\\ v(k,\mathbf{d}_{2};\theta_{q},\phi_{q})\\ \vdots\\ v(k,\mathbf{d}_{M};\theta_{q},\phi_{q})\end{bmatrix},\quad\forall q=1,...,Q,$ (2)
where $\mathbf{d}_{m}$ is the Cartesian coordinates of the $m$-th microphone in the array and $v(k,\mathbf{d}_{m};\theta_{q},\phi_{q})$ is the transfer function between a far-field source with a DOA of $(\theta_{q},\phi_{q})$ and $\mathbf{d}_{m}$ for $m=1,2,\ldots,M$. These steering vectors can be calculated analytically for various array types [17, 34], numerically, or measured.
@@ 2.2 Binaural Signal Representation using HRTFs
Assume that a listener is surrounded by a sound field that can be described by the plane-wave density (PWD) function $a(k,\theta,\phi)$. Then, the sound pressure at the listener’s ears can be described by [35]:
@@
$\mathbf{h}^{l,r}(k)=\begin{bmatrix}h^{l,r}(k,\theta_{1},\phi_{1})\\ h^{l,r}(k,\theta_{2},\phi_{2})\\ \vdots\\ h^{l,r}(k,\theta_{Q},\phi_{Q})\end{bmatrix}$
is a vector of length $Q$ containing the HRTFs corresponding to the $Q$ directions of the sources. These signals can be transformed to the time domain using the inverse discrete Fourier transform (IDFT) applied to (4) up to the Nyquist frequency. Then, they can be played back over headphones for binaural reproduction.
@@ 2.3 Beamforming-based Binaural Reproduction
where the beamformer output $y(k,\theta_{d},\phi_{d})$ constitutes the estimated PWD,
$\mathbf{w}_{d}(k)=\begin{bmatrix}w_{1}(k,\theta_{d},\phi_{d})\\ w_{2}(k,\theta_{d},\phi_{d})\\ \vdots\\ w_{M}(k,\theta_{d},\phi_{d})\end{bmatrix}$ (6)
@@
where $\{\alpha_{d}\}_{d=1}^{D}$ are the scaling factors. In BFBR design, the beamformer weights, the set of look-directions and the scaling factors are chosen with the aim of producing the desired binaural signals. The scaling factors provide additional control over the reproduced signal. For instance, certain look-directions that may be of high importance can be amplified by controlling the corresponding $\alpha_{d}$. In addition, with spherical arrays they can be set to produce binaural signals that are equivalent to HOA reproduction [23].
@@ 2.4 Binaural Reproduction using Ambisonics
where $\tilde{a}_{nm}(k)$ and $h_{nm}^{l,r}(k)$ are the spherical Fourier transform (SFT) coefficients of the complex conjugate of $a(k,\theta,\phi)$ and of $h^{l,r}(k,\theta,\phi)$, respectively, $N_{p}=\min\{N_{a},N_{H}\}$ and $(\cdot)^{*}$ is the complex conjugate operator. Equation (8) is using the Ambisonics signal, $a_{nm}(k)$, explicitly for binaural reproduction [2], and it is presented here since this formulation will be used in the listening experiment as a reference.
@@ 3 Proposed Method for Binaural Signal Matching
This section describes the proposed BSM method for binaural reproduction with general arrays. This method produces time-invariant multiple-input-multiple-output (MIMO) filters in the frequency domain. While the derivations of BSM shown here were already presented in previous works [24, 36, 13, 30], they are repeated here for completeness and to highlight the specific design parameters of BSM that are analyzed in the following section. These parameters are crucial for generating accurate binaural signals. For brevity, the wave-number index, $k$, will be omitted.
@@ 3.1 Formulation of the BSM Approach
$\hat{p}_{\text{BSM}}^{l,r}=[\mathbf{c}^{l,r}]^{H}\mathbf{x},$ (9)
where $\hat{p}_{\text{BSM}}^{l,r}$ are the estimated binaural signals according to the BSM method, and $\mathbf{c}^{l,r}$ are vectors of length $M$ holding the filter coefficients for the left and rights ears that will be formulated below. Next, the following MSE between the binaural signals in (4) and the filtered microphone signals in (9) is calculated for each ear separately:
$\epsilon^{l,r}=\mathbb{E}\begin{bmatrix}|p^{l,r}-\hat{p}_{\text{BSM}}^{l,r}|^{2}\end{bmatrix},$ (10)
where $\mathbb{E}\begin{bmatrix}\cdot\end{bmatrix}$ is the expectation operator. Assuming that the source signals, $\{s_{q}\}_{q=1}^{Q}$, are uncorrelated with the noise components, $\{n_{m}\}_{m=1}^{M}$, and substituting (1), (4) and (9) in (10), leads to the following MSE:
$\epsilon^{l,r}=$ $\mathbb{E}\left[|[\mathbf{h}^{l,r}]^{T}\mathbf{s}-[\mathbf{c}^{l,r}]^{H}\big{(}\mathbf{V}\mathbf{s}+\mathbf{n}\big{)}|^{2}\right]$
$=$ $\mathbb{E}\left[|\big{(}[\mathbf{h}^{l,r}]^{T}-[\mathbf{c}^{l,r}]^{H}\mathbf{V}\big{)}\mathbf{s}-[\mathbf{c}^{l,r}]^{H}\mathbf{n}\big{)}|^{2}\right]$
$=$ $\mathbb{E}\left[|\big{(}[\mathbf{c}^{l,r}]^{H}\mathbf{V}-[\mathbf{h}^{l,r}]^{T}\big{)}\mathbf{s}+[\mathbf{c}^{l,r}]^{H}\mathbf{n}\big{)}|^{2}\right]$
$=$ $\big{(}[\mathbf{c}^{l,r}]^{H}\mathbf{V}-[\mathbf{h}^{l,r}]^{T}\big{)}\mathbf{R_{s}}\big{(}[\mathbf{c}^{l,r}]^{H}\mathbf{V}-[\mathbf{h}^{l,r}]^{T}\big{)}^{H}$
@@
where, in the fourth line of (11) we have used the assumption that the source and noise signals are uncorrelated, and $\mathbf{R_{s}}=\mathbb{E}\left[\mathbf{ss}^{H}\right]$ and $\mathbf{R_{n}}=\mathbb{E}\begin{bmatrix}\mathbf{nn}^{H}\end{bmatrix}$ are the covariance matrices of the source and noise signals, respectively. When the sound sources and noise components are independent and identically distributed (i.i.d.), these matrices reduce to $\mathbf{R_{s}}=\sigma_{s}^{2}\mathbf{I}_{Q}$ and $\mathbf{R_{n}}=\sigma_{n}^{2}\mathbf{I}_{M}$, where $\sigma_{s}^{2}$ and $\sigma_{n}^{2}$ are the source and noise powers, respectively, and $\mathbf{I}_{Q}$ and $\mathbf{I}_{M}$ are the identity matrices of sizes $Q$ and $M$, respectively. In this case, (11) can be further simplified to
$\epsilon^{l,r}=$ $\sigma_{s}^{2}\left\lVert[\mathbf{c}^{l,r}]^{H}\mathbf{V}-[\mathbf{h}^{l,r}]^{T}\right\rVert_{2}^{2}+\sigma_{n}^{2}\left\lVert[\mathbf{c}^{l,r}]^{H}\right\rVert_{2}^{2}$
$=$ $\sigma_{s}^{2}\left\lVert\mathbf{V}^{H}\mathbf{c}^{l,r}-[\mathbf{h}^{l,r}]^{*}\right\rVert_{2}^{2}+\sigma_{n}^{2}\left\lVert\mathbf{c}^{l,r}\right\rVert_{2}^{2},$ (12)
where $\left\lVert\cdot\right\rVert_{2}$ is the $l^{2}$-norm, and the second term in (12) can be interpreted as Tikhonov-regularization [37]. The following derivations of BSM utilize the simplified model in (12), which requires less information on the sound field compared to the model in (11). While this simplified model may not always hold in practice, it will be shown later in Section 4.2 that using (12) may not limit the practical use of BSM, and that the formulations do hold for more general sound fields in (11).
@@
$\hat{p}_{\text{BSM}}^{l,r}=[\mathbf{c}^{l,r}_{\text{BSM}}]^{H}\mathbf{x}.$ (15)